Find the slope and the y-intercept of the line. Write fractional answers in the lowest terms. 6y+8x+5=20+8x. The slope is?
Answer by stanbon(75887)

Find the slope and the y-intercept of the line. Write fractional answers in the lowest terms. 6y+8x+5=20+8x. The slope is?
----------
Put in slope-intercept form:
6y = 15
y = 0x + (3/2)
---------------------
slope = 0
---
intercept = (3/2)
===================
Cheers,
Stan H.